Eight times the first of three consecutive odd integers is 4 less than  thrice the sum of second and third. The second integer is      ___

Grade:8

1 Answers

Vikas TU
3 years ago
Dear student 
Let, the three consecutive odd integers be (x+1) , (x+3) , (x+5)
 
So,
Eight times the first number = 8(x+1)
Sum of second and third number = (x+3) + (x+5)
 
According to Sum,
 
8 ( x + 1 ) + 4 = 3 [( x + 3 ) + ( x + 5 )]
↬8x + 8 + 4 = 3 [ x + 3 + x + 5 ]
↬ 8x + 12 = 3 ( 2x + 8)
↬ 8x + 12 = 6x + 24
↬ 8x - 6x = 24 - 12
↬ 2x = 12
↬ x = 6
 
Second number =
⇢ ( x + 3 )
⇢ ( 6 + 3 )
⇢ 9